OpenStudy (anonymous):

Check my answer please! Find the equation of the line in slope-intercept form that is perpendicular to y=-1/4x+3 and passes through (3/2,5) answer: y=-4x+3?

OpenStudy (mathstudent55):

Without any calculations, I see you a have an error.

OpenStudy (mathstudent55):

The slopes of perpendicular lines are negative reciprocals, not just reciprocals. That means the slopes multiply to -1. If one slope is - 1/4, the slope of a perpendicular must be 4, not -4.

OpenStudy (anonymous):

so it's just positive 4 right?

OpenStudy (mathstudent55):

The slope is positive 4, correct. You need to find the correct b.

OpenStudy (mathstudent55):

The perpendicular line has equation y = 4x + b Find the value of b.

OpenStudy (anonymous):

so y=4x+11?

OpenStudy (mathstudent55):

No.

OpenStudy (mathstudent55):

How did you get 11? Can you show me your work?

OpenStudy (anonymous):

wait is it -1?

OpenStudy (mathstudent55):

Correct. If you guess enough answers, sooner or later, you'll guess the correct one. In a test, though there won't be anyone you can ask all choices who will answer yes or no until you find the right one. I think it's a good idea to know how to solve it.

OpenStudy (anonymous):

This is my work. 5=4(3/2)+b 4x(3/2)=6 and then i forgot what i did after that THANK YOU! lol

OpenStudy (mathstudent55):

Very good! This is how it goes: 5=4(3/2)+b 5 = 12/2 + b 5 = 6 + b -1 = b b = -1 Now we insert -1 for b, and we get: y = 4x - 1 Good job! Sorry for assuming you were guessing.